OpenStudy (anonymous):

Solve by elimnation or linear combination- 5x-y=-8 & 5x + 4y =7

OpenStudy (anonymous):

sum both lines: - 5x-y=-8 (1) 5x + 4y =7 (2) ----------- 3y=-1 >>>>y=-1/3 y in (2) x=(7-4y)/5=5/3